13. What volume would 55.2 g acetylene gas, C2H2, occupy at a temperature of 25.0°C and a pressure of 1.25 atm?
Answer #1

From ideal gas equation, PV = nRT

Volume occupied by the acetylene gas, V = nRT/P

V = 55.2*0.0821*298/26*1.25 = 41.55 liters